# Sensitive 3mm Imaging of Discrete Sources in the Fields of tSZ-Selected Galaxy Clusters

### v1

@@ 5 Comparisons with other cluster samples
Figure 6: The average 90 GHz flux from point sources as a function of radius for this work (M2-ACT), the D2021 sample and the Websky simulation. For a fair comparison, the M2-ACT survey selection function has been applied to the other two samples. In the central $1^{\prime}$ the Websky simulation shows a slight excess of flux compared to the ACT sample however further out the ACT sample has a clear excess of flux from radio sources. The clear excess of sources in the D2021 sample can be seen in the central bin. For reference, the response to point sources of the ACT DR5 cluster finder is shown in grey.
@@
The D2021 clusters were all deep observations in order to study the physics of the intracluster medium or to measure the cluster masses. Many clusters in this non-uniform sample were selected for study as they were known to be mergers or at high redshifts and are atypical. The sample shows a much greater density of sources in the cluster centers (Figure 6). This persists even when the completeness limits of this work are taken into account. Interestingly, this is not the case outside the central ${}^{\prime}$, where the ACT clusters in this work show more radio emission from sources. Sources in this region increase $y_{c}$ and thus increase the likelihood of a cluster being detected by experiments such as ACT. Because of the unrepresentative nature and small size of the D2021 sample it is hard to draw more quantitative conclusions from this.
@@ 5.2 Comparisons with Websky
Websky is a set of simulated observations of the extragalactic microwave sky (Stein et al., 2020). It is used for forecasting for next generation CMB experiments (e.g., Hensley et al., 2022), evaluating biases in analysis methods (e.g., Coulton et al., 2023), and providing theory foregrounds (e.g., Madhavacheril et al., 2023; Qu et al., 2023). The Websky maps come with a corresponding catalog of radio sources (Li et al., 2022) and a map of their millimeter emission. This catalog is generated from the halos present in the base Websky observations, and whose distribution of intrinsic fluxes and spectral indices has been calibrated to match observed literature values in the millimeter. One of the primary uses of this radio catalog is to understand biases imparted in tSZ and lensing measurements by radio sources. While the radio fluxes do not bias the lensing reconstruction itself, when using lensing mass calibrations, clusters must be binned in mass to achieve sufficient signal-to-noise. This binning is done using the tSZ-inferred mass, and hence can be biased by radio source contamination (Madhavacheril et al., 2023). Therefore, confirming that the Websky radio accurately reproduces the observed tSZ contamination is of high priority.
To this end, we compare the results of the M2-ACT survey to the corresponding Websky simulations. Specifically, we select a $\sim$10000 cluster subsample of the Websky halos that matches the redshift and mass distributions of the ACT DR5 cluster catalog used in this work. We then identify all Websky radio sources lying within $5^{\prime}$ radially of the center of the cluster and with flux at $90$ GHz $\geq 0.5$ mJy, which should include all sources detectable in this work. As can be seen in Figure 6, once the survey completeness is taken into account, the Websky sample shows a reasonable match to the M2-ACT sample within the central $1^{\prime}$ but as with D2021, outside this radius there is a clear excess of emission from radio sources from the clusters in the M2-ACT sample.
@@
There are two potential explanations for this. First, the Websky simulation always places radio sources at the centers of the clusters and the only sources further out are those not associated with that cluster. In practice, many real clusters contain point sources away from their centers. This is especially true at higher redshifts where sources may still be infalling towards the cluster centers. For this to be the explanation, the Websky data should show an excess of flux in their centers equal to that missing further out. Although there is a slight excess of flux in the central ${}^{\prime}$ of the Websky clusters, there is not enough of an excess to explain the excess of sources found between ${}^{\prime}$ and ${}^{\prime}$ in the M2-ACT sample. If the M2-ACT sources are split into a higher and lower redshift groups (at $z=0.4$), then it becomes clear that the majority of the sources observed away from the cluster centers are at higher redshifts. The Websky simulations do not capture this effect well.
@@
The second explanation is due to how a source affects the measured Compton-$y$. For the ACT DR5 pipeline sources less than ${}^{\prime\prime}$ of the cluster center (the null in Figure 1) reduce the measured Compton-$y$ so some clusters with strong sources in the center will be missed by tSZ surveys such as ACT. Sources further out increase the measured Compton-$y$ and could scatter lower mass clusters into the tSZ selected cluster sample. Compared to the underling cluster population a tSZ-selected cluster sample can be expected to have an underrepresented number of clusters with bright sources in their centers and an over representation of those further out. A more quantitative analysis follows.
@@
5.2.1 Implications for the ACT Y-M relation
For each cluster in the Websky and ACT samples $\Delta y_{c}$ can be calculated using Equation 1 and the locations and fluxes of any sources. As the Websky sample contains many sources below the sensitivity of the ACT sample we also applied the completeness function described in Section 3 to the Websky sources and recalculated $\Delta y_{c}$. A comparison of these distributions is shown in Figure 7. Note that this comparison only includes Websky radio sources, and not dusty ones. However, Orlowski-Scherer et al. 2021 found that the $y_{c}$ contamination by infra-red sources of ACT clusters is negligible as compared to that from radio sources. A similar result was found by Bleem et al. (2023) in data from the South Pole Telescope.
First we examine the effects of sources in the centers of the clusters which reduce the measured $y_{c}$ and are responsible for the left side of Figure 7. Once the Websky sample is adjusted to account for the survey limits of our sample then there is very good agreement between the two when the measured $y_{c}$ is between 0.8 and 1.0 times the true value. The Websky sample also predicts that 1.5% of clusters should have a measured $y_{c}$ less than 0.75 times its true value. No ACT clusters with sources strong and close enough to their centers to produce this large of an effect were observed. If the Websky survey distribution of such strong central sources were correct this would imply tSZ surveys will miss 1.5% of clusters. Without more data from a non-tSZ selected cluster sample it is not possible to confirm this number. For this reason, follow up observations of an X-ray selected cluster sample is planned.
Without the flux cut, the Websky sample predicts that 5% of clusters should have sources that reduce $y_{c}$ by more than 5%. The difference between Websky predictions of $\Delta y_{c}$ with and without the flux cut represents the contribution of sources with fluxes just under 1 mJy. Although the ACT sample measured in this paper can not place good limits on fainter sources it shows a good match to brighter sources from Websky in the center of the clusters. Consequently, it is likely that 5% of clusters have measured $y_{c}$ values 5% or more lower than their true value.
At larger distances from the cluster center, sources increase the measured $y_{c}$ from its true value. In Figure 7 this shows up on the right. There is disagreement between the Websky prediction and our sample. Our sample shows that 3.3% of clusters have measured $y_{c}$ 10% or more greater than the true value and 2% have values 20% greater or more, while Websky gives numbers of 1.5% and 0.7% respectively. Some clusters in our sample had values of $y_{c}$ 50% high.
@@ 6 Conclusions
Our survey of 243 ACT clusters detected 24 sources. Even allowing for different survey depths this is significantly lower than expected when extrapolating from D2021. Most (75%) of these sources had low frequency radio counterparts but deep, high resolution follow-up at other tSZ bands would help confirm detections and spectral indices. Within the central ${}^{\prime}$ of the cluster there is a reasonable match to Websky for sources in the 1–5 mJy range (which reduce measured $y_{c}$ by up to 80%). Missing from our survey are clusters with strong sources in their centers – likely because such sources would lower $y_{c}$ enough to exclude clusters from our tSZ selected cluster sample. A complimentary survey of X-ray selected clusters is planned in order to place limits on the number of bright 90 GHz sources in the centers of clusters.
@@
At larger radii from the cluster centers, our survey found an excess of sources brighter than $\sim$1 mJy compared to Websky. The planned X-ray sample will distinguish if this is a limitation in the Websky model or a selection effect of the tSZ sample. Regardless of the cause, the Websky model does not reproduce the same positive changes in $y_{c}$. If a deeper MUSTANG2 survey had been possible it is likely the discrepancy with Websky would have become larger. Using a simple extrapolation to fainter sources based on the relative numbers of fainter Websky sources, we calculate that over 3% of clusters would have a 20% or more increase in measured $y_{c}$. In a stacking analysis of our results, sources in the outer regions increase the average mass bin by 1% (compared to 0.3% for the Websky prediction). While small this value could easily increase to 2% once fainter sources are included, a level which if unaccounted for, is comparable to the desired accuracy of future tSZ surveys. Once our X-ray survey is complete, comparison with other source models such as Sehgal (Sehgal et al., 2010) or Agora (Omori, 2022) simulations would be interesting.
@@
When working with individual cluster masses from tSZ surveys there is a chance of very significant ($>$30%) errors in the individual masses due to radio sources. Extrapolations from low frequency radio surveys do not predict this well. With observations often taken decades apart, source variability may play a role here but by far the biggest effect is the huge range in spectral indices of the sources. Confirmation of masses by other methods such as weak lensing or searches for sources with better than 1 mJy sensitivity at frequencies around tSZ wavelengths would be needed for science results that depend on the mass of an individual cluster.
@@
In summary, the Websky model for point sources is sufficient for today’s cluster surveys but in the future such models may need to include offsets of the sources from the tSZ centers. Predicting the distributions of sources from low frequency data is not possible. Instead, high resolution observations in the tSZ frequency bands are needed to confirm models for point source distributions in clusters. These observations should include non-tSZ cluster samples and measurements at multiple frequencies in order to better classify source properties.
@@ acknowledgments
Construction and early operations of MUSTANG2 was supported by the NSF award number 1615604 and the Mt. Cuba Astronomical Foundation. The GBT data were acquired under the project IDs AGBT21B_298, and AGBT22B_242. The Green Bank Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. SRD is supported by NSF grant number 2307546. MH acknowledges financial support from the National Research Foundation of South Africa. LDM acknowledges support from the French government, through the UCAJ.E.D.I. Investments in the Future project managed by the National Research Agency (ANR) with the reference number ANR-15-IDEX-01.

### v2

@@ 5 Comparisons with Other cluster samples
[…]
@@
The D2021 clusters were all deep observations in order to study the physics of the ICM or to measure the cluster masses. Many clusters in this nonuniform sample were selected for study because they were known to be mergers or at high redshifts and are atypical. The total fraction of clusters containing sources in the D2021 clusters is higher than that of the M2-ACT sample. However, the significance of this is highly dependent on assumptions such as the underlying distribution of source fluxes. Using only the total number of sources found, the hypothesis that the D2021 and the M2-ACT survey samples are drawn from the same parent sample of clusters can be rejected by as little as $1.2\sigma$ to well over $3\sigma$.
It is far more instructive to look at the fluxes and locations of the sources in each sample. The D2021 sample shows a much greater density of sources in the cluster centers (Figure 6). This persists even when the completeness limits of this work are taken into account. Interestingly, this is not the case outside the central 52′′, where the ACT clusters in this work show more radio emission from sources. Sources outside of $104^{\prime\prime}$ increase the measured Compton-$y$ and thus increase the likelihood of a cluster being detected by experiments such as ACT. The different distributions of sources within each cluster sample is further evidence that they are not drawn from the same parent cluster sample. Because of the unrepresentative nature and small size of the D2021 sample, it should not be used to correct $Y$-$M$ relationships.
@@ 5.2 Comparisons with Websky
Websky is a set of simulated observations of the extragalactic microwave sky (Stein et al., 2020). It is used for forecasting for next-generation CMB experiments (e.g., Hensley et al., 2022), evaluating biases in analysis methods (e.g., Coulton et al., 2023), and providing theory foregrounds (e.g., Madhavacheril et al., 2024; Qu et al., 2024). The Websky maps come with a corresponding catalog of radio sources (Li et al., 2022) and a map of their millimeter emission. This catalog is generated from the halos present in the base Websky observations and whose distribution of intrinsic fluxes and spectral indexes has been calibrated to match observed literature values in the millimeter. One of the primary uses of this radio catalog is to understand biases imparted in tSZ and lensing measurements by radio sources. While the radio fluxes do not bias the lensing reconstruction itself, when using lensing mass calibrations, clusters must be binned in mass to achieve sufficient SNR. This binning is done using the tSZ-inferred mass and hence can be biased by radio source contamination (Madhavacheril et al., 2024). Therefore, confirming that the Websky radio accurately reproduces the observed tSZ contamination is of high priority.
[…]
@@
There are two potential explanations for this. First, the Websky simulation always places radio sources at the centers of the clusters. The only sources further out are those not associated with that cluster. In practice, many real clusters contain point sources away from their centers. This is especially true at higher redshifts, where sources may still be infalling toward the cluster centers. For this to be the explanation, the Websky data should show an excess of flux in their centers equal to that missing further out. Although there is an excess of flux in the central $52^{\prime\prime}$ of the Websky clusters, there is not enough of an excess to explain the excess of sources found between 52′′ and 3 in the M2-ACT sample. If the M2-ACT sources are split into a higher- and lower-redshift groups (at $z=0.4$), then it becomes clear that the majority of the sources observed away from the cluster centers are at higher redshifts. The Websky simulations do not capture this change in source distribution with redshift well, but as the excess of flux in their centers is not equal to the deficit further out, the different distributions of sources within the clusters cannot be the only explanation.
@@
The second explanation is due to how a source affects the measured Compton-$y$. For the ACT DR5 pipeline, sources less than 104′′ from the cluster center (the null in Figure 1) reduce the measured Compton-$y$ so some clusters with strong sources in the center will be missed by tSZ surveys such as ACT. Sources further out increase the measured Compton-$y$ and could scatter lower-mass clusters into the tSZ-selected cluster sample. Compared to the underlying cluster population, a tSZ-selected cluster sample can be expected to have an underrepresented number of clusters with bright sources in their centers and an overrepresentation of those further out (past 104′′). A more quantitative analysis follows.
@@
6 Implications for the ACT $Y$-$M$ relation
For each cluster in the Websky and ACT samples, $\Delta y_{c}$ can be calculated using Equation 1 and the locations and fluxes of any sources. As the Websky sample contains many sources below the sensitivity of the ACT sample, we also applied the completeness function described in Section 3 to the Websky sources and recalculated $\Delta y_{c}$. A comparison of these distributions is shown in Figure 7. Note that this comparison only includes Websky radio sources, and not dusty ones. However, Orlowski-Scherer et al. 2021 found that contamination by infrared sources of ACT clusters is negligible as compared to that from radio sources. A similar result was found by Bleem et al. (2024) in data from the SPT.
First, we examine the effects of sources in the centers of the clusters that reduce the measured $y_{c}$ and are responsible for the left side of Figure 7. Once the Websky sample is adjusted to account for the survey limits of our sample, there is very good agreement between the two when the measured $y_{c}$ is between 0.8 and 1.0 times the true value. Although this may seem like a contradiction to the excess flux shown in the Websky sample between $0^{\prime\prime}$ and $52^{\prime\prime}$ in Figure 6, once the response to a point source and the excess flux in the M2-ACT sample between $52^{\prime\prime}$ and $104^{\prime\prime}$ is taken into account, the results are consistent. The Websky sample also predicts that 1.5% of the clusters should have a measured $y_{c}$ less than 0.75 times its true value. No ACT clusters with sources strong and close enough to their centers to produce this large of an effect were observed. If the Websky survey distribution of such strong central sources were correct, this would imply that tSZ surveys will miss 1.5% of clusters. Without more data from a non-tSZ selected cluster sample it is not possible to confirm this number. For this reason, follow-up observations of an X-ray-selected cluster sample are planned.
Without the flux cut, the Websky sample predicts that 5% of clusters should have sources that reduce $y_{c}$ by more than 5%. The difference between Websky predictions of $\Delta y_{c}$ with and without the flux cut represents the contribution of sources with fluxes just under 1 mJy. Although the ACT sample measured in this paper cannot place good limits on fainter sources, it shows a good match to brighter sources from Websky in the center of the clusters. Consequently, it is likely that 5% of clusters have measured $y_{c}$ values 5% or more lower than their true value.
At larger distances from the cluster center, sources increase the measured $y_{c}$ from its true value. In Figure 7 this shows up on the right. There is disagreement between the Websky prediction and our sample. Our sample shows that 3.3% of clusters have measured $y_{c}$ 10% or more greater than the true value and 2% have values 20% greater or more, while Websky gives numbers of 1.5% and 0.7%, respectively. Some clusters in our sample had values of $y_{c}$ 50% high.
@@ 7 Conclusions
Our survey of 243 ACT clusters detected 24 sources. Even allowing for different survey depths, this is significantly lower than expected when extrapolating from D2021. Most (75%) of these sources had low-frequency radio counterparts, but deep, high-resolution follow-up at other tSZ bands would help confirm detections and spectral indexes. Within the central $104^{\prime\prime}$ of the clusters, there is a reasonable agreement on the effect on $y_{c}$ with Websky for sources in the 1–5 mJy range (which reduce the measured $y_{c}$ by up to 80%). Missing from our survey are clusters with strong sources in their centers, likely because such sources would lower $y_{c}$ enough to exclude clusters from our tSZ-selected cluster sample. A complimentary survey of X-ray-selected clusters is planned in order to place limits on the number of bright 90 GHz sources in the centers of clusters.
@@
At larger radii from the cluster centers, our survey found an excess of sources brighter than $\sim$1 mJy compared to Websky. The planned X-ray sample will distinguish if this is a limitation in the Websky model or a selection effect of the tSZ sample. Regardless of the cause, the Websky model does not reproduce the same positive changes in $y_{c}$. If a deeper MUSTANG2 survey had been possible, it is likely the discrepancy with Websky would have become larger. Using a simple extrapolation to fainter sources based on the relative numbers of fainter Websky sources, we calculate that over 3% of clusters would have a 20% or more increase in measured $y_{c}$. In a stacking analysis of our results, sources in the outer regions increase the average mass bin by 1% (compared to 0.3% for the Websky prediction). While small, this value could easily increase to 2% once fainter sources are included, a level that if unaccounted for, is comparable to the desired accuracy of future tSZ surveys. Once our X-ray survey is complete, comparison with other source models such as Sehgal (Sehgal et al., 2010) or Agora (Omori, 2024) simulations would be interesting.
@@
When working with individual cluster masses from tSZ surveys, there is a chance of very significant ($>$30%) errors in the individual masses due to radio sources. Extrapolations from low-frequency radio surveys do not predict this well. With observations often taken decades apart, source variability may play a role here, but by far the biggest effect is the huge range in spectral indexes of the sources. Confirmation of masses by other methods such as weak lensing or searches for sources with better than 1 mJy sensitivity at frequencies around tSZ wavelengths would be needed for science results that depend on the mass of an individual cluster.
@@
In summary, the Websky model for point sources is sufficient for today’s cluster surveys, but in the future, such models may need to include offsets of the sources from the tSZ centers. Predicting the distributions of sources from low-frequency data is not possible. Instead, high-resolution observations in the tSZ frequency bands are needed to confirm models for point-source distributions in clusters. These observations should include non-tSZ cluster samples and measurements at multiple frequencies in order to better classify source properties.
@@ acknowledgments
Construction and early operations of MUSTANG2 were supported by NSF award No. 1615604 and the Mt. Cuba Astronomical Foundation. The GBT data were acquired under the project IDs AGBT21B_298, and AGBT22B_242. The Green Bank Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This publication makes use of data products from the Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. S.R.D. is supported by NSF grant No. 2307546. M.M. acknowledges support from NSF grants AST-2307727 and AST-2153201 and NASA grant 21-ATP21-0145. M.H. acknowledges financial support from the National Research Foundation of South Africa. L.D.M. acknowledges support from the French government, through the UCAJ.E.D.I. Investments in the Future project managed by the National Research Agency (ANR) with the reference No. ANR-15-IDEX-01.